FarCry Industries, a maker of telecommunications equipment, has 5 million shares of common stock outstanding, 2 million shares of preferred stock outstanding, and 20,000 bonds. Suppose the common shares are selling for $29 per share, the preferred shares are selling for $15.50 per share, and the bonds are selling for 98 percent of par.

What would be the weight used for equity in the computation of FarCry’s WACC? (Round your answer to 2 decimal places.)
Answer #1

Total value of

Common stock=(5*29)=$145 million

Preferred stock=(2*15.5)=$31 million

Bonds=20,000*(1000*98%)=$19.6 million

Total value=(145+31+19.6)=195.6 million

Hence weight of equity=(145/195.6)

=74.13%(Approx).